OpenStudy (izzy18970):

I really need help solving this !!!! A gym sells memberships for $325. To determine the amount of money made on memberships, the gym can use the equation, t=325m, where t represents the total amount of money made on memberships and m represents the amount of memberships sold. If the gym sells 19 memberships in January, how much money will be made on memberships for the month? $ If the gym sells the same amount of memberships each month, how much money will be made in a year? $

OpenStudy (anonymous):

So to solve this equation, you can substitute the number of memberships sold into the equation. t= 325(m) m=19 t=325(19) multiply 325*19 to get 6175 $6175 will be made on memberships for the month. There are 12 months in the year. If we multiply the number of months by the amount of money made each month, we can find the total amount of money made in the year. 6175*12 = 74100 $74 100 will be made on memberships for the year. Does this help?

OpenStudy (anonymous):

do you go to connections?

OpenStudy (anonymous):

Nope!

OpenStudy (izzy18970):

that's what I got. It just looked wrong